a thin circular wire of radius r has a charge Q. if a point charge q is placed at the centre of the ring, then find the increase in tension in the wire.

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
At the point when the charge q=+Q is given to the wire then it comes in pressure which is called circle strain. At the point when at the focal point of the ring another charge q0=+Q'is given, the pressure in the wire increments to keep this fragment into equillibrium once more. 
Presently we can consider the basic strip subtended at an edge Ѳ at its inside. 
As the string is in balance in x bearing. For harmony in y course. 
F = 2T sin(dѳ/2)......................... (1.) 
q/2π xdѳ 
Charge on the strip 
F = 1/4 πɛ x q/2π x(dѳ xqo)/r2 - - (2.) 
Subsitute 2 of every 1 
2T dѳ/2 =1/4 πɛ x q/2π x(dѳ xqo)/r2 
T = qq0/8π2ɛr2.
 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)